News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 15 August 2008 9.00am EDT

On 12 Aug Chandra experienced what is believed to be a single event upset (SEU) in the telemetry processor of the Command and Telemetry Unit, causing a reset of its timing signals. The Observatory was slewing between targets at the time of the presumed SEU. The timing reset slightly modified the end-point of the maneuver, which resulted in the spacecraft entering normal-sun mode. As a result of the timing reset, several subsystem hardware items also experienced a reset.

Recovery procedures were successfully executed on 14 Aug to recover to normal-point mode. Prior to running the procedure the ephemeris-time-offset K-constant was patched in the flight software. This patch was necessary for the on-board computer to be able to calculate the position of the Sun for solar array pointing. Chandra was also maneuvered to a cooler attitude as part of the procedure.

Following the recovery from normal-sun mode a clean-up procedure was run to re-load software patches to the EPHIN Input/Output unit. Clean-up tasks for the aspect camera electronics, science instrument module electronics, and ACIS had been performed earlier.

Science will be resumed this morning at 8:52am.

The schedule of targets following the resumption of science and for the next week is shown below.

------------------------------------------
PSR B1957+20           ACIS-S       Aug 16
Radiation Belts                     Aug 18
Carina Clusters Pt     ACIS-I       Aug 19
2MASSW J1139511-31     ACIS-S
Carina Clusters Pt     ACIS-I       Aug 20
cl2220-5228            ACIS-I
J0940-5428             ACIS-I
NGC 6681               ACIS-S       Aug 21
Radiation Belts
WR 140                 ACIS-S/HETG
MACSJ1931.8-2634       ACIS-I       Aug 22
IGR J13135-1109        ACIS-I       Aug 23
J1841-0345             ACIS-I
NGC 6638               ACIS-S
Radiation Belts
WR 140                 ACIS-S/HETG  Aug 24
M81 DwA                ACIS-S
Holmberg IX            ACIS-S       Aug 25

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports